LHC Tech Stop July 2011 Action plan for

  • Slides: 6
Download presentation
LHC Tech Stop July 2011 Action plan for major stops (+multi instrum & other

LHC Tech Stop July 2011 Action plan for major stops (+multi instrum & other works) SC - 08 Jul'11 1. 8 K 70 mba r RF RF Powerin g Quench p SAM um g. P Ref. CP 7 Ref. De 1. 8 K DFB g. P u ? mp SAM 15 mba r. DFB 15 mba ? Test ? r repump <50 mba r Ve um p L Je De g. P Empty 1. 8 K QRL_ C QR SAM DFB Me Re fill Ma De 12 23 34 45 56 67 78 81 Lu 70 mba r 15 mba DFB r LHCCryo - Tech. Stop July 2011

And reality … Major finding: the screw compressor (P 8, oil filters clogging mid

And reality … Major finding: the screw compressor (P 8, oil filters clogging mid June) was found damaged, with bearings showing exceptional stress About 35’ 000 hrs, bearings calculated for 70’ 000 hrs Similar case than P 4_Xmas: - Series problem ? - Too much axial force ? Confirmed by observations on oil circuit => Decision to treat similar compressors SC - 08 Jul'11 a. s. a. p Outlet (20 bar) Inlet (4 bar) Specific oil injection (22 bar) to balance axial forces, Need to send more oil than pass through to screws ! LHCCryo - Tech. Stop July 2011

Compressor station of LHC 18 k. W@ 4. 5 K Pipe & fittings redone

Compressor station of LHC 18 k. [email protected] 4. 5 K Pipe & fittings redone with larger diameters SC - 08 Jul'11 Bldg: 15 x 45 X 9 Pinput : 4. 5 MW Cool: 500 m 3/h Noise: 105 d. BA P 8(1) done this Tech Stop, 2 nd to factory, P 4(1) & P 18(2) done in shadow of P 8 recovery, LHCCryo - Tech. Stop July 2011 P 6 to be done latest at Next Tech Stop

P 8 configuration, week-end Extra flow Validated for P 2 (TS) ! Tests for

P 8 configuration, week-end Extra flow Validated for P 2 (TS) ! Tests for relocation Until compressor repaired! Limitations Leaky HX (Jan’ 11) Tu. A (Spring 11) 6 CV 215 valve (P 2) CCs QUI Sector 78 SC - 08 Jul'11 Some minimum flexibility recovered during Tech Stop Sector 81 LHCCryo - Tech. Stop July 2011

LHC Tech Stop July 2011 Tentative d’illustration des activités Cryo LHC 1. 8 K

LHC Tech Stop July 2011 Tentative d’illustration des activités Cryo LHC 1. 8 K um p QR Re fill p SAM um De Ref. CP 7 Ref. g. P u ? mp 1. 8 K DFB 70 mba r Ve SAM 15 mba 15 CP r. DFBCP mbar 15 mba ? Test r ? Re-pump repump CP 15 mba <50 mba r r Powerin g Quench 70 mbar UPS SAM 70 mbar 70 mba r w Ne 1. 8 K Je L Me D SAM Empty 1. 8 K QRL_ FB C De 12 23 34 45 56 67 78 81 Ma De g. P Lu 15 mba DFB r RF RF Powerin g Quench Rev 2: P 5_UPS effect + refinement 12/34 SC - 08 Jul'11 Rev 3: QRL_12 test stopped earlier + P 4/P 18 LHCCryo - Tech. Stop July 2011 QSCB

Status & Forecasts Overall situation under control, checking singularities here and there P 8:

Status & Forecasts Overall situation under control, checking singularities here and there P 8: Should be back Friday evening (22 h t. b. c) Special: ALL triplets + D 4 Q 5_R 4 emptied, warmed-up 25 K, pumped VAC SC - 08 Jul'11 LHCCryo - Tech. Stop July 2011